Department Fabricating Machining Assembly Total Plant Manufacturing overhead Direct labor 379,750 434,000 97,650 $ 911,4co s 217,000 108,s00 325,500 651,000 Jobs require varying amounts of work in the three departments. he Koopers job,for example, would have required manu costs in the three departments as follows: Direct materials Direct labor Manufacturing overhead Fabricating 4,700 6,200 hining 300 600 Assembly $3,100 7,900 Total Plant 8,100 $14,700 Required: 1. Using the companys plantwide approach: a. Compute the plantwide predetermined rate for the current yeat b. Determine the amount of manufacturing overhead cost that would have been applied to the Koopers job. 2. Suppose that instead of using a plantwide predetermined overhead rate, the company had used departmental predetermined overhead rates based on direct labor cost. Under these conditions: a. Compute the predetermined overhead rate for each department for the current year b. Determine the amount of manufacturing overhead cost that would have been applied to the Koopers job. 4. Assume that it is customary in the industry to bid obs at 150% of total manufacturing cost (direct materials, direct labor and applied overhead). a. What was the companys bid price on the Koopers job using a plantwide predetermined overhead rate? b. What would the bid price have been if departmental predetermined overhead rates had been used to apply overhead cost?
Answer #1

Answers

  • All working forms part of the answer
  • Requirement 1

---Part ‘a’

A

Total Manufacturing Overhead

$911,400

B

Total Direct Labor Cost

$651,000

C = A/B

Plantwide predetermined rate

$                     1.40 or 140% = Answer

--Part ‘b’

A

Total Direct Labor cost

$14,700

B

Plantwide predetermined rate

$                     1.40 or 140%

C = A x B

Amount of manufacturing Overhead allocated

$20,580 = Answer

  • Requirement 2

--Part ‘a’

Working

Fabricating

Machining

Assembly

A

Manufacturing Overhead

$379,750

$434,000

$97,650

B

Direct Labor cost

$217,000

$108,500

$325,500

C = A/B

Departmental rates = ANSWER

$                     1.75 or 175%

$                    4.00 or 400%

$                   0.30 or 30%

--Part ‘b’

Working

Fabricating

Machining

Assembly

TOTAL

A

Direct Labor

$6,200

$600

$7,900

$14,700

B

Departmental rates

$                     1.75

$                    4.00

$                   0.30

C = A x B

Manufacturing Overhead applied = Answer

$10,850

$2,400

$2,370

$15,620 = ANSWER

  • Requirement 4

--Working and answers

Working

Requirement 4'a'

Requirement 4'b'

A

Direct material cost

$8,100

$8,100

B

Direct Labor cost

$14,700

$14,700

C

Manufacturing Overhead applied

$20,580

$15,620

D = A+B+C

Total cost

$43,380

$38,420

E

Bid price % on total cost

150%

150%

F = D x E

Bid Price [ANSWERS]

$65,070 = Answer

$57,630 = Answer